# Neural-Driven Image Editing

### v1

@@ A.2 Subject information
12 healthy college students (6 females and 6 males) were recruited as the subjects for data collection. They have a mean age of 24.5$\pm$2.5 years and normal (or corrected-to-normal) vision. All volunteers were informed of the experimental process and received financial compensation. All volunteers signed the consent forms prior to the experiment, which was approved by the Ethics Committee of the corresponding Institutional Review Board for Human Research Protections (Protocol ID: No. 067 (2019)). The attention score in this study is objectively computed using the ratio between the power of the EEG alpha band (8–12 Hz) and the theta band (4–8 Hz), as shown below:
Attention Score=Alpha Band PowerTheta Band $\text{Attention Score}=\frac{\text{Alpha Band Power}}{\text{Theta Band Power}}$ (13)
This ratio has been widely used in cognitive neuroscience research as a neurophysiological index of attentional control and mental workload. For instance, Raufi [64] demonstrated that the alpha-to-theta and theta-to-alpha band ratios are reliable indicators of self-reported mental workload levels in EEG-based studies.
Table 4: Summary of attention scores and neural-signal based image editing metrics per subject

### v2

@@ A.2 Subject information
[…]
$\text{Attention Score}=\frac{\text{Alpha Band Power}}{\text{Theta Band Power}}$ (13)
This ratio has been widely used in cognitive neuroscience research as a neurophysiological index of attentional control and mental workload. For instance, Raufi [64] demonstrated that the alpha-to-theta and theta-to-alpha band ratios are reliable indicators of self-reported mental workload levels in EEG-based studies. The attention scores and the neural-signal based image editing errors of each individual subject are summarized in Table.4.
Table 4: Attention and neural-signal based image editing scores of individual subject